# Learning A Physical-aware Diffusion Model Based on Transformer for Underwater Image Enhancement

### v1

@@ 5 Conclusion
In this paper, we develop a novel UIE framework, namely PA-Diff. With utilizing physics prior information to guide the diffusion process, PA-Diff can obtain underwater-aware ability and model the complex distribution in real-world underwater scenes. To the best of our knowledge, it is the first the Diffusion model based on physical perception in image enhancement tasks. Our designed physics prior generation branch is a plug-and-play universal module to produce physics prior information, which can be integrated into any deep learning framework. PA-Diff shows SOTA performance on UIE task, and extensive ablation experiments can prove that each of our contributions is effective.

### v2

@@ 5. Conclusion
In this paper, we develop a novel UIE framework, namely PA-Diff. With utilizing physics prior information to guide the diffusion process, PA-Diff can obtain underwater-aware ability and model the complex distribution in real-world underwater scenes. To the best of our knowledge, it is the first the diffusion model based on physical perception in image enhancement tasks. Our designed PPU is a plug-and-play universal module to exploit the beneficial feature-level physics information, which can be integrated into any deep learning network. PA-Diff shows SOTA performance on UIE task, and extensive ablation experiments can prove that each of our contributions is effective.